T : R

3 → R

2

 defined by : 9 

T(x, y, z) = (x -y + z, -2x + 2y -2 z)


1
Expert's answer
2021-09-07T16:43:04-0400

1. Let "u=(u_1, u_2, u_3)" and "v=(v_1, v_2, v_3)" be vectors in "R^3" and "c" and "d" be scalars.

Consider

"T(cu+dv)=T(cu_1+dv_1, cu_2+dv_2, cu_3+dv_3)"

"=(cu_1+dv_1-(cu_2+dv_2)+cu_3+dv_3,"

"-2(cu_1+dv_1)+2(cu_2+dv_2)-2(cu_3+dv_3))"

"=(c(u_1-u_2+u_3)+d(v_1-v_2+v_3),"

"c(-2u_1+2u_2-2u_3)+d(-2v_1+2v_2-2v_3))"


"=c(u_1-u_2+u_3, -2u_1+2u_2-2u_3)"

"+d(v_1-v_2+v_3, -2v_1+2v_2-2v_3)"

"=cT(u)+dT(v)"

Therefore the transformation "T:R^3\to R^2," given by

"T(x, y, z)=(x-y+z, -2x+2y-2z)"

is linear.


2.


"T(e_1)=T(1, 0,0)"

"=(1-0+0, -2+0-0)=(1, -2)"


"T(e_2)=T(0, 1,0)"

"=(0-1+0, -0+2-0)=(-1, 2)"


"T(e_3)=T(0, 0,1)"

"=(0-0+1, -0+0-2)=(1, -2)"

So


"A=\begin{pmatrix}
 1 & -1 & 1 \\
 -2 & 2 & -2
\end{pmatrix}"

3.

"T(0,0,0)=(0,0)"